OpenStudy (mollykenney):

A convex lens has a diameter of 30 mm and a depth of 5 mm at the center. How far from the vertex is the focus?

Elsa213 (elsa213):

@ganeshie8 might help c:

OpenStudy (mollykenney):

i hope so @Elsa213 !

OpenStudy (michele_laino):

here the focal distance \(f\) is: \[\frac{1}{f} = \frac{2}{R}\] wherein \(R\) is the radius of the lens, namely: \[R = \frac{D}{2} = \frac{{30}}{2} = ...mm\] so the requested distance \(d\), is: \(d=f-(5/2)=...\)